A projectile is fired from the surface of the earth with a speed of 200m/s at an angle of 300 above the horizontal. If the ground is level, what is the maximum height reached by the projectile


1
Expert's answer
2021-03-31T09:39:14-0400

Let's consider "g \approx 10 \; m/s^2".

The maximum height in projectile motions is

"\displaystyle h_{max}=\frac{v_0^2 \sin^2 \alpha}{2g} = \frac{200^2 \cdot 0.5^2 }{2 \cdot 10} = \frac{100^2}{20} = 500 \,m"


Answer: 500 m